OpenStudy (anonymous):

Simplify : 1/4 (1•2)+ -1 Help please! Thanks in advanced.

OpenStudy (krabsen):

1/4(1x2)+(-1)= 1/4x2/4+(-1)= 2/4+(-1)= 1/2+(-1)= -1/2

OpenStudy (krabsen):

hope that helps